OpenStudy (anonymous):

5/8 divided by 4/3 perform the indicated operation.

OpenStudy (anonymous):

Remember that you can always multiply any number by any factor that equals 1. In this case, you could multiply by (3/4)/(3/4) to eliminate the denominator, as ((3/4)/((3/4)*(4/3)) = 3/4. This gives you (5/8)*(3/4) = 15/32